INAF queue observing, 2018 Jan 27-28

We (the observers) take this space to bring attention of the world to the the awkward situation of large part of research personnel in INAF. We are now performing observations for INAF researchers with LBT, one of the biggest telescope in the world. But we are also valued for our scientific contribution, for results that we publish as first authors and for those in which we are involved.

However, we believe that like for many other INAF post-docs, our contract situation and our expectations do not reflect the great results that we, and INAF thanks to us, are obtaing. Like many other post-docs, we are paid on contracts renewed every year pending the availability of external funds.

Moreover, we are not entitled to officially say a word on institutional choice and in many aspects we are not considered as INAF employees, but like external collaborators. Indeed, we cannot even lead INAF projects and build a independent research.

Now, we believe INAF has given the power and instruments to start changing this situation.

Observer: R. Carini, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O. Kuhn
Instrument: LUCI2

Summary

At the sunset the sky is clear instruments: LBC works, LUCI1 only imaging and AO, LUCI2 imaging and LS and MOS, MODS1 works, MODS2 only blue channel

Night duration (12\xB0 el twilight) 2:00am- 13:15pm UT = 11 h 15m

Observing time: 11 h 15m (100%)

Weather time loss: 0 h (0%)

technical time loss 0 h (0%)

Data Summary

Project_RUN Objects Instrument Filter ExpTime Notes
Nisini telluric field 3 HIP 17971 LUCI2 J,K,H    
Nisini Field3 LUCI2 J 15 m filter J done
Vanzella JDz11 LUCI1-AO Ks 30 m   24 m in spec
telluric J1 HIP30393 LUCI2 zjspec, HKspec    
Antoniucci J1 agll 192 LUCI2 Hkspec, zjspec 0.5 h spectra
telluric I1 HIP30393 LUCI2 zjpec, HKspec    
Antoniucci K2 Dr Tau LUCI2 HKspec, zjspec 0.4h +0.1 spec +ima completed
Antoniucci C4 Xz tau LUCI2 HKspec, zjspec 0.4h to be repeated
telluric K2 HIP34612 LUCI2 HKspec, zjspec    
Antoniucci K2, K3 IPTF15AFQ LUCI2 Hkspec, zjspec 0.3 miss 2 images in HKspec and ima
telluric Mignoli 45666 LUCI2 HKspec    
Mignoli B 401129 LUCI2 HKspec 2h  
telluric Mignoli HIP 54815 LUCI2 HKspec(lambda=1.93)    
telluric Piconcelli A5 LUCI2 LUCI2 HKspec (lambda=1.95)    
Piconcelli A5 J1110+1930 LUCI2 HKspec   COMPLETED
Arc+Flat Nisini Field3 LUCI2 H,J,K    
Dark Vanzella        
flat Vanzella   Ks  

Details

01:37 begin with Nisini program, sand the ob of the telluric for the Field3
  • acq luci2.20180128.0006-0009
  • sci luci2.0010-0021
  • seeing 0.9''

01:57 slew to the scientific field3
  • acq luci2.0021-
  • seeing 0.5" on guider
  • failed guiding , because too close to zenith. We wait a bit
  • no good mask allignement, we try again without allignement star 30
  • 2:42 sill alligning . Finally alll almost inside, impossible to center all of them!
  • 2:50 sci luci2.0030-0032 J band completed, y=0 and Y=+2 offsets

03:10 slew to AO, Vanzella program
  • 03:30 facing the wind, but let's go, keep setting up collimation and AO
  • 03:35 error with the rotator, TO sends again the preset
  • 03:38 send the script
  • sci luci1.20180128.0006-0020
  • seeing 0.9''on the guider
  • 03:41 error again
  • send the script again, luci1_SL_Vanz.xml
  • seeing 0.7'' on the guider
  • measured magnitude on the reference AO star is 13.4 mag
  • first image luci1.0006 FWHM=0.25''
  • luci1.0007 FWHM=0.22''
  • luci1.0008 FWHM=0.22''
  • luci1.0010 FWHM=0.21''
  • ==luci1.0013 ==FWHM=0.22''
  • ==luci1.0020 ==FWHM=0.22''
  • luci1 have taken just 14 images instead 26, we don't know why!!!! The last with big offset are missing !!!
  • 04:10 the measured magnitude of the reference AO star is 13.6 because clouds come.

04:13 send the script again luci1_SL_Vanz.xml
  • sci ==luci1.20180128.0021-0028==FWHM=0.22''
  • 04:40 ==luci1.0029-0031==not good, cirrus passing by in the last 10 minutes

04:49 the seeing is too high and the clouds too thick, we stop AO
  • we try to take the imaging seeing limited for the calibration of Vanzella, but the script give us error in LUCI1 (script said to move the mask but it is impossible because it has been disabled this option ), LUCI1 is stuck
Later , Olga find oput that Andrea made a mistake in the script, not putting N30 filed stop mask in FPU frown, sad smile

04:57 slew to Antoniucci program, telluric star, seeing 1.6''
  • the TO have to collimate again with LUCI2
  • in the meantime Dave Thompson resolves the problem with LUCI1, now it works.
  • *05:09*send the script agll_192_tell.xml
  • acq luci2.0033-0035 seeing variable 1.0''-1.8''
  • sci luci2.0036-0039 seeing 1.3''-1.4''

05:22 slew to the the science target AGLL192
  • acq luci2.0040-0043 seeing 1.3''-1.4''
  • 05:38 do acquisition again, because wrong offset.
  • seeing 1.5" on dimm and guider at 5:40
  • acq luci2.0044-0047
  • sci luci2.0048-0055 seeing 1.6"
  • 05:50 seeing 1.4"

06:15 the seeing is bad and variable, we stay on Antoniucci Program, telluric for AGLL 354:
  • acq luci2.0056-0058
  • sci luci2.0059-0062
  • seeing between 1.0''and 1.4''

06:33 send the scientific script for AGLL 354
  • error during the preset
  • there are problems with the guide star, maybe in the script

06:43 send a script for another target V512 Per, again error during the preset

06:45 we slew to Dr Tau
  • acq luci2.0063-0066
  • sci spectra luci2.0067-0074
  • seeing 1.00''-1.50''
  • sci ima luci2.0075-0080

07:09 slew to XZ Tau
  • acq luci2.0081-0084
  • sci spec luci2.0085-0092 seeing variable between 1.2'' to 1.8''
  • the script has the wrong central wavelength for the HKspec filter, so these images are not good. The target is too low and the TO said us to slew to another one.

07:34 slew to the tellurica star HIP34612 for the target IPTF15AFQ
  • acq luci2.0093-0095 seeing 1.5''-2.0''
  • sci luci2.0096-0099 seeing 1.5''-1.6''

07:46 slew to the target IPTF15AFQ
  • acq luci2.0100-0103
  • sci spec 0104-0109 during the integration of the second image in the HKspec the guide star has been lost, but the spectrum is there.

08:39 the seeing seems stable, less than 1.2'', we slew to Mignoli program, we start with the telluric 45666
  • acq luci2.0110-0112- seeing about 1.2''
  • sci luci2.0113-0116 seeing about 1.2'' on the DIMM

08:56 slew to the target 401129
  • acq luci2.0117-0166 seeing variable between 0.7''and 1.3'' we put the reference star at about y=1688.
  • sci luci2.0167-0184 the first image (0167) has to be excluded because it has readout lir. we have changed in mer.
  • 09:32 seeing between 0.80''-1.00''
  • 09:41 seeing 0.8" . We can see the continuum of target and ref star in the spectra. We are very good
  • 09:52 seeing 0.7''
  • 10:02 seeing 0.6''
  • 10:26 seeing variable 0.85''-1.2''

10:34 seeing variable 0.8''-1.10'', there are thin clouds., we send again the script for MIgnoli program
  • sci luci2.0185-0203
  • 10:58 seen 0.8''-1.0''
  • 11:37 seeing 0.8''-1.3''
  • we take one more image to finish the 2h. seeing 0.9''

11:56 othe telluric for Mignoli
  • acq luci2.0204-206
  • sci luci2.207-210 seeing about 1.2''

12:11 telluric for Picocnelli A5
  • acq luci2.0211-0213
  • sci luci2.0214-0215

12:22 Target Piconcelli A5
  • faild the preset, TO cheks pointing
  • 12:26 send the script again
  • acq luci2.0216-0219
  • sci luci2.0220-0231
  • seeing 1.3''

* The calibration for Nisini taken just now are not good for the data taken today because the Field stop has been just aligned, we do them again for the Field 3.
  • flat in H luci2.0237-0241
  • arc in H luci2.0242-0243 Argon
  • arc in H luci2.0244-0245 Xenon
  • flat in J luci2.0251-0255
  • arc J luci2.0256-0257 Argon
  • arc J ==luci2.0256-0257==Xenon
  • flat K luci2.0260-0269
  • arc K ==luci2.0270-0271==Argon
  • arc K luci2.0272-0273 Xenon
  • arc K luci2.0274-0275 Neon

* dark for Vanzella luci1.0032-0036 * flat Ks for Vanzella luci1.0037-0041

* Calib for Antoniucci:
  • arc HK ==luci2.0276-0278==Neon arg xe, NOT OK because no dark in between

CIAO

-- %USERSIG{RobertaCarini - 2018-01-28}%

Comments

 
Topic revision: r21 - 30 Jan 2018, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback